# Diffraction of large-number whispering gallery mode by boundary straightening with jump of curvature

### v1

@@ 5.2 Investigation of integral $I_{2}$
[…]
@@ 6 Conclusions
Diffraction of a large-number whispering gallery mode by jumply straitening of curvature has been considered. The wavefield has been described in detail in a small neighbourhood (see (15)) of the jump point $O$ to the right of it within the framework of the parabolic equation method. A comprehensive geometrical analysis of the problem has been carried out, which allows to give a geometrical interpretation to the derived asymptotic formulas. Let us summarise results obtained above and compare them with the ones of the study [12] devoted to the case of a small-number mode.
@@
For large $\mathfrak{t}$, terms (68), (69), (84) corresponding to the rays are clearly distinguishable in the asymptotics of the field. [^5] Note that contributions of the rays close to the horizontal ray $l_{B}$ (see Fig. 5), as mentioned in section 5.1 and especially around (70), can not be described analytically. Apparently, this is because of the insufficient accuracy of approximation (9) for the incident wave. To the right of $O$, a rather extended caustic emerges. In its vicinity (blue zone in Fig. 5) the wavefield is described by the sum of expression (80) involving the Airy function and diffracted wave (96)–(100). Noteworthy is that diffraction coefficient (97)–(100) is not linear in the amplitude of the curvature jump $1/a$, whereas for $\mathfrak{t}=O(1)$ and the Neumann boundary condition it is.
In the case of moderate $\mathfrak{t}$, the diffracted wave has a singularity on the straight part of the boundary. In its vicinity, the wavefield is described by the parabolic cylinder functions $D_{-4}$ with arguments independent of the geometrical parameters of the problem, see [12]. However, in the case of large $\mathfrak{t}$, diffracted wave (96)–(100) is singular on the limit ray $l_{O}$, but not on $S_{+}$, see Fig. 5. Transition zone surrounding $l_{O}$ is divided into two parts by the point $Q$. To the left of $Q$ (green zone in Fig. 5), the wavefield is expressed by the sum of the Fresnel integrals (85) and (74) and the regular part of diffracted wave (see (96), (97)). Formulas (85) and (74) describe the merging of the singular part of diffracted wave (see (96), (98)–(100)) with the wave $u_{1}$ after its reflection from $S_{+}$ and with the wave $u_{2}$ before passing the caustic, respectively. To the right of $Q$ (yellow zone in Fig. 5), the asymptotics of wavefield is the sum of the Fresnel integrals (85) and (71) and the regular part of diffracted wave. Formula (71) describes the merging of the singular part of diffracted wave with the wave $u_{2}$ after passing the caustic and has an imaginary argument, which is unusual for diffraction problems. Expressions (86), (75) and (73) in the arguments of the Fresnel integrals depend on the curvature radius of $S_{-}$ and the grazing angle $\gamma$ (see Fig. 2).
Finally, in the boundary layer surrounding the point $Q$ (red zone in Fig. 5) where the caustic touches the limit ray, the wavefield is described by the sum of incomplete Airy function (82) and the regular part of diffracted wave.

### v2

@@ 5.2 Investigation of integral $I_{2}$
$I_{2}^{\pm}=\frac{e^{-i\pi/4}}{2\sqrt{\pi\sigma}}\int_{L_{2}}v(p-\mathfrak{t})e^{i\frac{(p\pm\nu)^{2}}{4\sigma}}dp.$ (101)
@@ 6 Conclusions
Diffraction of a large-number whispering gallery mode by jumply straitening of curvature has been considered. The wavefield has been described in detail in a small neighborhood (see (16)) of the jump point $O$ to the right of it within the framework of the parabolic equation method. A comprehensive geometrical analysis of the problem has been carried out, which allows to give a geometrical interpretation to the derived asymptotic formulas. Let us summarise results obtained above and compare them with the ones of the study [12] devoted to the case of a small-number mode.
@@
For large $\mathfrak{t}$, terms (69), (70), (85) corresponding to the rays are clearly distinguishable in the asymptotics of the field. [^5] Note that contributions of the rays close to the horizontal ray $l_{B}$ (see Fig. 5), as mentioned in section 5.1 and especially around (71), can not be described analytically. Apparently, this is because of the insufficient accuracy of approximation (9) for the incident wave. To the right of $O$, a rather extended caustic emerges. In its vicinity (blue zone in Fig. 5) the wavefield is described by the sum of expression (81) involving the Airy function and diffracted wave (2), (98). Noteworthy is that in the area where $\varphi\gg\gamma$ diffraction coefficient (98) coincides with the one for $\mathfrak{t}=O(1)$. Moreover, formula (98) agrees with the expression for diffraction coefficient for non-tangential incidence.
In the case of moderate $\mathfrak{t}$, the diffracted wave has a singularity on the straight part of the boundary. In its vicinity, the wavefield is described by the parabolic cylinder functions $D_{-4}$ with arguments independent of the geometrical parameters of the problem, see [12]. However, in the case of large $\mathfrak{t}$, diffracted wave (2), (98) is singular on the limit ray $l_{O}$, but not on $S_{+}$, see Fig. 5. Transition zone surrounding $l_{O}$ is divided into two parts by the point $Q$. To the left of $Q$ (green zone in Fig. 5), the wavefield is expressed by sum of the Fresnel integrals (86) and (75) and the background (88), (91). Formulas (86) and (75) describe merging of diffracted wave with the wave $u_{1}$ after its reflection from $S_{+}$ and with the wave $u_{2}$ before passing the caustic, respectively. To the right of $Q$ (yellow zone in Fig. 5), the asymptotics of wavefield is sum of the Fresnel integrals (86) and (72) and the background. Formula (72) describes merging of diffracted wave with the wave $u_{2}$ after passing the caustic and has an imaginary argument, which is unusual for diffraction problems. Expressions (87), (76) and (74) in the arguments of the Fresnel integrals depend on curvature radius of $S_{-}$ and grazing angle $\gamma$ (see Fig. 2).
Finally, in the boundary layer surrounding the point $Q$ (red zone in Fig. 5) where the caustic touches the limit ray, the wavefield is described by the sum of incomplete Airy function (83) and the background.